In how many ways we can select 5 cards from a deck from 52 cards, if each selection must include atleast one king?

Grade:11

1 Answers

Naveen Kumar
9 years ago
As there are 4 cardd of king in the set of 52 cards.
total ways of selection of 5 cards from the deck = 52C5
case-I. If we select all the 4 king in the selection so the number of ways=52-4=48=l(say)
case-II. If only one king then total ways=48C4=(m say)
case-III.similarly for selection of 2 kings, total ways=48C3 =n(say)
case-IV.for selection of 3 kings, total ways=48C2=t(say)
So total number of ways=l+m+n+t